OpenStudy (megannicole51):

Find the sum of the series... -2, 1-(1/2)+(1/4)-(1/8)+(1/16) so i know A1=-2 and r=(-1/2) thanks to @hartnn ....and i was going to use the formula A1=(1-r^n/1-r) but im not sure what n is:(

OpenStudy (megannicole51):

is there another formula if you dont know what n is?

hartnn (hartnn):

yes, correct r = -1/2 just plug in that formula n is just the number of terms, you can count it manually,

OpenStudy (megannicole51):

so n=6?

hartnn (hartnn):

yup

OpenStudy (megannicole51):

awesome:) that was easy lol

hartnn (hartnn):

unless there is a ..... in the end

hartnn (hartnn):

which would make it infinite term series

hartnn (hartnn):

-2, 1-(1/2)+(1/4)-(1/8)+(1/16) ...... <<<<<infinite series sum = a/(1-r)

OpenStudy (megannicole51):

i dont see an a so im good with the first formula right? lol

hartnn (hartnn):

the formula is \(\huge S_n= a_1\dfrac{1-r^n}{1-r}\) where a1 is the first term

hartnn (hartnn):

just to confirm .....as i saw A1=... in your question...

OpenStudy (megannicole51):

perfect thats what i used:) i got -21/16

hartnn (hartnn):

i get that too! :)

OpenStudy (megannicole51):

yay:) i have another question ill repost it:) hopefully it will be as easy as this one!